OpenStudy (anonymous):

simplify square root of 72x^3y^4

OpenStudy (anonymous):

how does 3x*9y*\[\sqrt{8x}\]sound

OpenStudy (anonymous):

\[6\sqrt{2} x^{3/2}y^{2}\]

OpenStudy (anonymous):

6xy^2 root2x

OpenStudy (anonymous):

bracket can you show how you did that please, we need to show the work

OpenStudy (anonymous):

root (a*b*c) = root(a) * root(b) * root(c) so\[ \sqrt{72x ^{3}}y ^{4} = \sqrt{72}*\sqrt{x ^{3}}*\sqrt{y ^{4}}\]

OpenStudy (anonymous):

and 72=2*2*2*3*3=2^2*3^2*2

OpenStudy (anonymous):

\[\sqrt{x ^{3}} = (x ^{3})^{1/2} = x ^{3/2}\]

OpenStudy (anonymous):

\[\sqrt{y ^{4}} = (y ^{4})^{1/2} = y ^{4/2} = y ^{2}\]

OpenStudy (anonymous):

so\[\sqrt{72}x ^{3}y ^{4}=6\sqrt{2}*x ^{3/2}*y ^{2}\]

OpenStudy (anonymous):

Hope this clear??? If yes, pls click the GOOD ANSWER button for me :)